COMPARING THE ORTHO-TO-PARA RATIOS OF H_{2} AND H_{3}^{+} IN DIFFUSE INTERSTELLAR CLOUDS

\underline{NICK INDRIOLO}, Department of Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801; TAKESHI OKA, Department of Astronomy & Astrophysics and Department of Chemistry, University of Chicago, Chicago, IL 60637; THOMAS R. GEBALLE, Gemini Observatory, Hilo, HI 96720; KENNETH H. HINKLE, National Optical Astronomy Observatories, Tucson, AZ 85726; GEOFFREY A. BLAKE, Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

The ratio between the populations of the two lowest rotational levels of H_{2}, J = 0 and J = 1, can be used to determine the temperature of interstellar gas (referred to as T_{01}). Likewise, a temperature can be inferred from the populations of the (J, K) = (1, 0) and (J, K) = (1, 1) states of H_{3}^{+}. However, the average temperatures derived from these methods (T_{01} ≈ 60 K, T (H_{3}^{+}) ≈ 30 K) do not agree. Theories predict that the deviation from a Boltzmann distribution in both species is due to collisions between H_{2} and H_{3}^{+} which can change the spin alignment. Recent laboratory results confirm this deviation from a thermal distribution, and provide a relationship between the (1,0)/(1,1) ratio of H_{3}^{+} and the (1)/(0) ratio of H_{2}. We have made observations searching for H_{3}^{+} in several sight lines with measured H_{2} abundances for the purpose of determining this relationship in interstellar clouds. With such a relationship, we then show that IR observations probing the (1,0) and (1,1) states of H_{3}^{+} can be used to estimate the H_{2} temperature in highly extincted sight lines where UV spectroscopy is not possible.